# Adaptive Active Inference Agents for Heterogeneous and Lifelong Federated Learning thanks: *Corresponding author

### v1

@@ V-A3 Application Scenario
We emulate a dynamic environment by controlling the data generation process to introduce concept and volume drifts. Each client device represents a different participant. The challenge is, for the system to adapt to the drifts or to the varying computational resources of participants. The experiments consider the fulfillment of two high-level SLOs:
@@
We choose time and performance as SLOs as balancing them is non-trivial. For example, focusing exclusively on fulfilling prediction performance may require spending an excessive amount of time and vice versa. The configurable hyperparameters are Batch Size $BS\in\{8,32,64,256,512\}$ and Learning Rate $LR\in\{0.0005,0.001,0.005,0.01\}$, as there is a clear connection to them and the system’s training objective and considered drift types, e.g. learning rate tuning was proposed to battle concept drift [11]. Each client initialized a separate datastream locally, using their identifier as a random seed for the stream to promote variability in the data across clients. However, the rest of the data stream parameters were set to the same values across clients inside one run. This way, the number of features and classes are kept the same, showcasing that all clients are part of the same ML task. Other parameters, such as the number of clusters (groups of data points represented by centroids), the random seed for cluster initialization, and the cluster drift, are kept constant to ensure a proper comparison of the agents’/clients’ behavior and to maintain the reproducibility of the generated data inside one repetition. The data generator $G$ is also parametrized by a drift parameter, $drift$, which controls the presence and speed of data drifts, where $drift=0$ indicates no data drift.
@@ V-A4 Baselines
The baselines focus on the best- and worst-case scenarios rather than existing methods to avoid misleading comparisons with the tangentially related problem definitions and methods and concentrate on depicting the dynamics and capabilities of AIF agents. We define two baselines for experiments with data distribution drifts:
@@
Each experiment was repeated ten times with different random seeds that controlled the random processes, such as the data generator and ANN weights initialization. For the evaluation, the first considered timestep included in SLO fulfillment tracking for a particular client run was the one where the “lifelong” flag becomes true. The reported results are averaged across all participating clients and experiments, if not stated otherwise. The number of local epochs was set to 3 for all experiments with no client subsampling, i.e., all FL clients participated in every round. The SGD optimizer was used by default. It is also worth reminding that agents learn their BNs from scratch in every experiment.
V-B Results
V-B1 Data Heterogeneity
Consider a real-life situation where clients experience changes in the amount of collected data, e.g., seasonal demands in shops specializing in certain types of products or bursts of the number of service requests. For this experiment, such quantity drift is modeled by increasing the number of samples drawn from the data generator every 50 epochs, starting with 5,000 samples, then increasing to 10,000 and 15,000. The SLOs were set to 2 seconds for time and 97% for model performance SLO. The number of federated clients was set to 10.
The dynamic of both SLO fulfillments is shown in Figure 4.
@@ V-B1 Data Heterogeneity
The EFE shows that having not challenging or unrealistic SLOs (Figure 8 a and b, respectively) leads to EFE being consistent across all available configurations, with EFE being high when the target SLOs are unfulfillable and EFE being equally low when the targets are too easy to fulfill. The situation is different when one of the two SLOs is relaxed while the other is somehow challenging to fulfill. Thus, when performance is relaxed, the agents are incentivized to optimize the behavior towards time SLO, leading to agents settling for the largest available batch size (512) regardless of the learning rate. The situation is different regarding time-relaxed SLO – when the performance is targeted, agents tend to explore more configurations as the task is still not that challenging for them, leading to diverse behaviors. Still, the resulting EFE heatmap expresses some preference bias compared to the experiment with two SLOs relaxed. This “uncertain” behavior could be attributed to quantity drifts not impacting the performance goals. A situation where time is relaxed in the presence of concept drift is presented later in this section, illustrating the preferences developed by the agents when model performance (and performance SLO) has a clearer connection with the configurations available.
The following experiment added concept drift to the problem. Concept drift was added to the generator from the start to make the task more complicated, and, as a result, the performance SLO was adjusted to still be challenging yet reachable (85%). One quantity drift was also present (from 10,000 to 15,000 samples), therefore, introducing two sources of data heterogeneity in one experiment. The resulting SLOs fulfillment for 5 FL clients is shown in Figure 9.
@@
Figure 9: Mean SLOs fulfillment under concept drift and 1 quantity drift. Time SLO: 3 seconds, performance SLO: 85%
With concept drift, reaching the target performance SLO becomes more challenging. Moreover, when the quantity drift happens – concept drift becomes more prominent in response, as more samples are drawn from the data generator and clusters drift faster. It impacts the model’s performance, as can be seen from the prominent decline of the random baseline, illustrating that fewer configurations can now sustain the performance goals. Despite that, the AIF agent maintains its performance SLO fulfillment and keeps time SLO not impacted by the quantity drift. Here, the agents preferred configurations with larger batch sizes of 256 and 512, with (512 0,0005) having the lowest mean EFE (compared to the fixed baseline being (256, 0.005)).
As was previously noted and could be seen from the previous example, concept drift mostly impacts the performance SLO, which is a more nuanced target compared to the time SLO, which is more straightforward to optimize. The experiment containing only more severe concept drift was conducted to better inspect the behavior of AIF agents when facing performance-oriented SLO. Here, the performance SLO was set to 78.5% (due to a more prominent concept drift) and the time SLO was relaxed (3600 s). The performance SLO fulfillment is shown in Figure 10 and the resulting mean EFE is shown in Figure 11.
Figure 10: Mean SLOs fulfillment under concept drift and relaxed time SLO. Time SLO: 2 seconds, performance SLO: 78.5%
Figure 11: Mean EFE under a concept drift with relaxed time SLO
The optimal configuration for the baseline was set to (64, 0.001), and it is shown from the mean EFE that agents also preferred a lower batch size of 32 and lower learning rates of 0.0005 and 0.001.
V-B2 Device Heterogeneity
@@
The next experiment focused on inspecting the ability of AIF agents to adapt to the resources available to the agents located at the edge devices and aligning them with the SLOs set. For the experiments, three edge devices (Raspberry Pi 5, Nvidia Orin NX, and Nvidia Xavier NX) were used as separate federated clients and were tasked to participate in the FL for 75 epochs, while a Raspberry Pi 4 device served as the FL orchestrator. No data drifts were introduced in this experiment. Due to the differences in resources, it was expected that agents would prefer different configurations under the same target SLOs. The comparison of SLO fulfillment across various devices is shown in Figure 12, and mean EFE is shown in Figure 13.
@@
It is shown that all devices manage to fulfill the set SLOs. It is also worth noting that Raspberry Pi 5 managed to fulfill performance SLO slightly faster than devices that used GPU but occasionally struggled to maintain flawless time SLO fulfillment. When looking at mean EFE, it is clear that in order to better utilize its cores (as there is no GPU), Raspberry Pi can successfully employ a vast range of configurations due to the small size of the model (64 and 32 units in 2 layers), while devices with GPU choose bigger batch sizes to better utilize their parallelization capabilities.
To further inspect the adaptability of agents to different combinations of tasks, available resources, and set SLOs, the experiment was conducted with the same set of devices but with a wider neural network (5120 and 512 units compared to 64 and 32 used in the previous experiment). The time SLO was adjusted to 15 seconds. This change was introduced to give GPU-empowered devices an advantage over Raspberry Pi 5. The resulting SLOs fulfillment is shown in Figure 14 and mean EFE is shown in Figure 15. The performance SLO fulfillment at the final FL round was 96.9% for Orin NX and 98.7% for RPI 5 and Xavier NX, while the time SLO was 99% for Nvidia devices and 87.8% for RPI 5.
@@
VI Related Work
We identified three topics within related work: SLO fulfillment through dynamic adaptations in distributed computing systems, applications of the AIF framework, and optimization techniques in FL.
VI-A Optimization in Federated Learning
Optimizing FL workflows is important for minimizing the time-to-accuracy of model training [23]. In that context, Kundroo et al. [9] proposed FedHPO, a federated optimization algorithm that accelerates each client’s training by modifying its hyperparameters, such as learning rate or epochs.
Furthermore, several studies have focused on multi-objective optimization in FL. One possible approach is optimizing neural network models instead of client-specific hyperparameter optimization [24, 25]. Additionally, a significant number of existing research focuses on optimizing client selection or clustering instead of adjusting the parameters of individual clients to reach multi-dimensional goals [26, 27, 28].
VI-B Dynamic Adaptation for SLO Fulfillment
There exist multiple approaches that aim to combine SLOs with dynamic processing requirements: Zhang et al. [29] presented Octopus – the framework that finds optimal services configurations in multi-tenant edge computing scenarios. Octopus predicts SLO fulfillment of two variables based on a deep neural network. Shubha et al. [8] presented AdaInf, which detects SLO violations of a GPU scheduling task whenever variable drifts occur. Through AdaInf, it is possible to find SLO-fulfilling resource allocations between model training and inference.
VI-C AIF Applications
@@ VII Conclusion
This work presented AIF agents that are able to adaptively change their behavior in response to dynamic environments. We evaluated the proposed AIF agents in lifelong heterogeneous FL, utilizing a set of both dynamic data and diverse devices. We showed that AIF agents are able to fulfill competing SLOs and unfolded the behaviors of agents and intricate connections between the defined SLOs and strategies for fulfilling them.
Future work can further expand the usage of the active inference framework to orchestrate distributed learning systems, for instance, by fulfilling system-level SLOs, such as fairness of participation or global model performance. Enhancements of the current method can improve the ability of the agents to find causal dependencies in limited data, making them more robust, targeting the limitations of the discrete BN, introducing temporal dependencies to capture the environmental dynamics more precisely, and providing more nuanced SLOs specifications to enable tracking SLOs in a range.

### v2

@@ V-A3 Application Scenario
We emulate a dynamic environment by controlling the data generation process to introduce concept and volume drifts. Each client device represents a different participant. The challenge is, for the system to adapt to the drifts or to the varying computational resources of participants. The experiments consider the fulfillment of two binary high-level SLOs:
@@
We choose time and performance as SLOs as balancing them is non-trivial. For example, focusing exclusively on fulfilling prediction performance may require spending an excessive amount of time and vice versa. The configurable hyperparameters are Batch Size $BS\in\{8,32,64,256,512\}$ and Learning Rate $LR\in\{0.0005,0.001,0.005,0.01\}$, as there is a clear connection to them and the system’s training objective and considered drift types, e.g., learning rate tuning was proposed to battle concept drift [12]. The values chosen were selected to span across reasonable ranges that can make a difference in terms of SLOs while being distinct from each other. Each client initializes an independent data stream locally.
The data generator $G$ is also parametrized by a drift parameter, $drift$, which controls the presence and speed of data drifts, where $drift=0$ indicates no data drift.
@@ V-A4 Baselines
We choose two baselines to present AIF agents. The first set focuses on the presentation of the behaviour of AIF agents, aiming at identifying both adaptable behaviour pattern and assessing agents’ performance. The second baseline compares AIF agents to the state-of-the-art Optuna framework to assess the optimality of hyperparameter choices made by the agents.
We define two baselines to illustrate AIF agents behaviour under data distribution drifts:
@@
To align Optuna with the AIF agent, its hyperparameter search process was modified. First, the hyperparameter study was set to cover all FL rounds, with the first hyperparameters set trial starting after the global model performance reached sufficient accuracy (same as for AIF agent as described in Section IV) and performing one trial per FL round until the end of the training. To select the configuration for the current FL round, first, an Optuna trial is performed and added to the study (stored in the local Optuna database). Then, the best configuration is inferred, and the FL training round starts. To maintain the integrity of Optuna trials in the lifelong learning process, each trial is conducted using a copy of the global model from the previous round, evaluated on the training data from that same round. By isolating trials from ongoing learning dynamics, we prevent information leakage in both directions – from the current global model to the trials and vice versa – preserving the validity of the optimization process.
To make Optuna SLO-aware, the study was designed in MOO fashion, where SLOs variables served as objectives. Here, training time and 1 – validation set accuracy served as two objectives for optimization (minimization). By default, for MOO Optuna returns not a single best configuration, but a Pareto front – a set of feasible configurations, leaving the final choice up to the user. In the experiments, we first attempt to filter out only those configurations in the Pareto front that fulfill both SLOs, then both time and performance metrics are normalized and weighted equally into one feature, where the best configuration is returned for the FL round training.
@@
Each experiment was repeated ten times with different random seeds that controlled the random processes, such as the data generator and ANN weights initialization. For the evaluation, the first considered timestep included in SLO fulfillment tracking for a particular client run was the one where the “lifelong” flag (model performance stabilizes and the agent starts learning) becomes true. The reported results are averaged across all participating clients and experiments, if not stated otherwise. The number of local epochs was set to 3 for all experiments with no client subsampling, i.e., all FL clients participated in every round. The SGD optimizer was used by default. It is also worth reminding that agents learn their BNs from scratch in every experiment.
V-B Agent Demonstration
Consider a real-life situation where clients experience changes in the amount of collected data, e.g., seasonal demands in shops specializing in certain types of products or bursts of the number of service requests. For this experiment, such quantity drift is modeled by increasing the number of samples drawn from the data generator every 50 epochs, starting with 5,000 samples, then increasing to 10,000 and 15,000. The SLOs were set to 2 seconds for time and 97% for model performance SLO. The number of federated clients was set to 10. The dynamic of both SLO fulfillments is shown in Figure 4.
@@ V-B Agent Demonstration
The EFE shows that having not challenging or unrealistic SLOs (Figure 8 a and b, respectively) leads to EFE being consistent across all available configurations, with EFE being high when the target SLOs are unfulfillable and EFE being equally low when the targets are too easy to fulfill. The situation is different when one of the two SLOs is relaxed while the other is somehow challenging to fulfill. Thus, when performance is relaxed, the agents are incentivized to optimize the behavior towards time SLO, leading to agents settling for the largest available batch size (512) regardless of the learning rate. The situation is different regarding time-relaxed SLO – when the performance is targeted, agents tend to explore more configurations as the task is still not that challenging for them, leading to diverse behaviors. Still, the resulting EFE heatmap expresses some preference bias compared to the experiment with two SLOs relaxed. This “uncertain” behavior could be attributed to quantity drifts not impacting the performance goals.
After dissecting the behaviour of the agents in a simpler scenario, the next section focuses on more practical experimental setups and presents evaluation of AIF agents’ behaviour in the presence of concept drifts and resources heterogeneity.
VI Results
VI-A Hyperparameter Tuning Comparison
To demonstrate both the validity and advantages of the proposed AIF agents, we compare them against Optuna, a state-of-the-art framework for hyperparameter optimization in ML. Optuna serves as a strong baseline due to its efficiency in finding optimal configurations. However, unlike Optuna, which focuses on static optimization, AIF agents offer continuous adaptability, making them better suited for dynamic and heterogeneous FL environments.
VI-A1 Concept Drift
The first experiment included a concept drift, which was present from the beginning of the training, and both approaches needed to find the best configuration to meet the time SLO of 3s and performance SLO of 85%. The SLOS fulfillment is shown in Figure 9.
@@ VI-A1 Concept Drift
Figure 9: Mean cumulative SLOs fulfillment for the scenario with a persistent concept drift
It is seen that both Optuna and AIF have similar time SLO fulfillment, with performance SLO being slightly better for AIF after training round 60.
To assess what configurations were favored by both approaches, EFE (Figure 10) and normalized objectives can be examined for AIF and Optuna, respectively.
Figure 10: Mean EFE for the scenario with a persistent concept drift
For AIF, the best configuration based on EFE is (64, 0.005). For Optuna, the result is based on separate time and performance assessments and their combination (Figure 11). For the time SLO, the best configurations are those of larger batch sizes, as expected. For performance, Optuna focused on a set of configurations from batch size 8 through 64 and learning rates of 0.001 through 0.01. The combination of both SLOs lands in the range of batch sizes (32, 64, 256) and learning rates of (0.005, 0.01), with (64, 0.005) being the most preferred configuration.
Figure 11: Optuna preferred configurations at the end (last 10 rounds) of training
Comparing the two approaches, it is seen that both AIF agents and Optuna converge to similar configurations in this scenario.
VI-A2 Concept and Quantity Drifts after Round 50
Next evaluation scenario introduced two drifts at the same time, simulating a change in both quantity and distribution of data received by the FL clients.
The scenario was set as follows. First, 10,000 samples were used per FL training round with no concept drift; after round 50 – a slight drift appeared and 20,000 samples started to be drawn from the data generator. The resulting SLOs fulfillment is presented in Figure 12 for time SLO – 4 seconds and performance SLO – 90%.
It is seen that adding concept drift to data leads to an immediate decline in performance SLO as previously established configurations and the trained model do not account for that. However, AIF agents are able to return to a relatively high performance SLO fulfillment in a faster and more reliable way than Optuna.
Figure 12: Mean cumulative SLOs fulfillment for the scenario with concept and quantity drifts after round 50
Based on EFE shown in Figure 13, it is seen that after the drifts appeared after round 50, a set of the previously favored configurations (256, 0.005) and (64, 0.005) gradually lost their importance and agents focused mostly on (256, 0.001).
Figure 13: Mean EFE for the scenario with concept and quantity drifts after round 50
These experiments illustrate that AIF agents align with commonly used hyperparameter optimization methods while providing adaptability where necessary. Compared to the state-of-the-art Optuna framework adapted to the lifelong heterogeneous FL scenario, AIF agents are able to fully recover from the concept drift after 43 rounds, while Optuna fails to return to its best performance even after 100 rounds after the drift introduction.
VI-A3 Device Heterogeneity
The next experiment focused on inspecting the ability of AIF agents to adapt to the resources available to the agents located at the edge devices and aligning them with the SLOs set.
@@ VI-A3 Device Heterogeneity
For the experiments, three edge devices (Raspberry Pi 5, Nvidia Orin NX, and Nvidia Xavier NX) were used as separate federated clients and were tasked to participate in the FL for 75 epochs, while a Raspberry Pi 4 device served as the FL orchestrator. No data drifts were introduced in this experiment. Due to the differences in resources, it was expected that agents would prefer different configurations under the same target SLOs. In addition to different available resources, the size of the neural networks was also changed, so to change the utilization of available resources on the device.
The first experiment featured a smaller neural network (64 and 32 units in 2 layers). The comparison of SLO fulfillment across various devices is shown in Figure 14, and mean EFE is shown in Figure 15.
@@
It is shown that all devices manage to fulfill the set SLOs. It is also worth noting that Raspberry Pi 5 managed to fulfill performance SLO slightly faster than devices that used GPU but occasionally struggled to maintain flawless time SLO fulfillment. When looking at mean EFE, it is clear that in order to better utilize its cores (as there is no GPU), Raspberry Pi can successfully employ a vast range of configurations, while devices with GPU choose bigger batch sizes to better utilize their parallelization capabilities.
The next experiment was conducted with the same set of devices but with a wider neural network (5120 and 512 units compared to 64 and 32 used in the previous experiment). The time SLO was adjusted to 15 seconds. The resulting SLOs fulfillment is shown in Figure 16 and mean EFE is shown in Figure 17. The performance SLO fulfillment at the final FL round was 96.9% for Orin NX and 98.7% for RPI 5 and Xavier NX, while the time SLO was 99% for Nvidia devices and 87.8% for RPI 5.
@@
Results presented in this section show how AIF agents display adaptive behaviour in the presence of changing data as well as adaptation to the local resources to ensure the fulfillment of set SLOs.
@@ VII Conclusion
This work presented AIF agents that are able to adaptively change their behavior in response to dynamic environments. We evaluated the proposed AIF agents in lifelong heterogeneous FL, utilizing a set of both dynamic data and diverse devices. We showed that AIF agents are able to fulfill competing SLOs and unfolded the behaviors of agents. We compared AIF agents to the hyperparameter-tuning framework Optuna adjusted to lifelong learning and showcased how the adaptive nature of AIF agents allows for faster performance recovery in the presence of data drift.
Future work can further expand the usage of the active inference framework to orchestrate distributed learning systems, for instance, by fulfilling system-level SLOs, such as fairness of participation or global model performance. Another line of research can target the scalability of the proposed framework, as the usage of BNs can potentially introduce computational bottleneck as the number of considered vertices and their cardinalities grow. Enhancements of the current method can improve the ability of the agents to find causal dependencies in limited data, making them more robust, targeting the limitations of the discrete BN, introducing temporal dependencies to capture the environmental dynamics more precisely, and providing more nuanced SLOs specifications to enable tracking SLOs in a range.